OpenStudy (anonymous):

simplify the trigonometric expression. sin^2(theta) / 1- cos(theta) A. 1+cos(theta) B. sin(theta) C. 1-sin(theta)/cos(theta) D. 1+sin(theta)/cos)theta)

OpenStudy (reemii):

you know \(\sin^2a+\cos^2 a = 1 \) ? afterwards use (a-b)(a+b)=a^2-b^2 ...

OpenStudy (anonymous):

i don't know how to do this at all..

OpenStudy (reemii):

the numerator is \(\sin^2\theta\). with the formula "that you should know", you can write instead \(1-\cos^2\). do that and then look at it again, thinking about how factorize it. (with the second formula i wrote)

OpenStudy (reemii):

since \(1-\cos^2\theta\) is a difference of squares (isn't it?)

OpenStudy (jhannybean):

\[\frac{\sin^2(\theta)}{(1-\cos(\theta)}\]\[\frac{(1-\cos^2(\theta))}{(1-\cos(\theta))}= \frac{(1+\cos(\theta))(1-\cos(\theta))}{(1-\cos(\theta))}\]\[(1+\cos(\theta))\]

OpenStudy (mathstudent55):